1.a) Calculate the molar mass, in g mol-1, of the following compounds, give your answers to 1dp:
i) Methane, CH4
ii) Calcium carbonate, CaCO3
iii) Potassium permanganate, KMnO4
iv) Ammonium sulphate, (NH4)2SO4
v) Iron(III)nitrate, Fe(NO3)3

b) Calculate the number of moles in the following, give your answers to 2 sig figs:

i) 23g of Na
ii) 36g of H2O
iii) 304g of Cr2O3 (chromium (III)oxide)
iv) 10g of CO2
v) 82g of C2H5OH (ethanol)
c) Calculate the mass, in g, of the following:
i) 2 moles of NH3 (ammonia)
ii) 0.5 moles of H2SO4 (sulphuric acid)
iii) 1.5 moles of NaCl (sodium chloride)
iv) 15 moles of Li2O (lithium oxide)
v) 3.45 moles of Sr(NO3)2 (strontium nitrate)
1
Expert's answer
2020-02-07T05:28:26-0500


  1. Calculate the molar mass as follows:

i) Mr (CH4)=Ar(C)+4Ar(H)=12+4*1=16 g/mol

ii) Mr (CaCO3)=Ar(Ca)+Ar(C)+3Ar(O)= 40+12+3*16=40+12+48=100 g/mol

iii)Mr(KMnO4)=Ar(K)+Ar(Mn)+4Ar(O)= 39+55+4*16=39+55+64= 158 g/mol

iv)М((NH4)2SO4) = 2*Ar(N) + 4*2*Ar(H) + Ar(S) + 4*Ar(O) = 2*14 + 8*1 + 32 + 4*16 =28+8+32+64=132 g/mol

v)М(Fe(NO3)3) = Ar(Fe) + 3*Ar(N) + 3*3*Ar(O) =56+3*14+9*16=56+42+144= 242 g/mol


2.Calculate the number of moles as follows:

v=m/Mr

i) v(Na)=23g/23g/mol=1mol

ii)v(H20)=36g/(2*1+16)g/mol=36 g/18g/mol=2mol

iii)v(Cr2O3)=304g/(2*52+3*16)g/mol=304g/152g/mol=2mol

iv)v(CO2)=10g/(12+2*16)g/mol=10g/44g/mol=0.2mol

v)v(C2H5OH)=82g/(2*12+5*1+16+1)g/mol=82g/46g/mol=1.8mol


3.Calculate the mass as follows:

m=v*Mr

i)m(NH3)=2mol*(14+3*1)g/mol=2*17=34 g

ii)m(H2SO4)=0.5mol*(2*1+32+4*16)g/mol=2*98=196 g

iii)m(NaCl)=1.5mol*(23+35)g/mol=1.5*58=87 g

iv)m(Li2O)=15mol*(2*7+16)g/mol=15*30=450 g

v)m(Sr(NO3)2) =3.45*(88+2*14+2*3*16)g/mol=3.45*212=731.4 g